alones (alones):

DID I get it right ;3

alones (alones):

(x^3 + 2x^2 + 2x + 1) ÷ (x + 1)

alones (alones):

x2 + x + 1 My steps were

OpenStudy (mathmale):

Are you allowed to use synthetic division here? If so, please do. Otherwise, do long division, showing all work. Note that x2 is ambiguous; please write that as x^2 or as \[x^2.\]

alones (alones):

Yeah i drew out my long division on a paper actually it will take a long time to draw ;3

alones (alones):

z^5+z^4+z^3+21z^2+9z+27z+3 =(z+3)(z^4−2z^3+7z^2+9)/z+3 =z^4−2z^3+7z^2+9

OpenStudy (alivejeremy):

No its totally wrong

alones (alones):

DAAMN!!!! My third day and i still have it wrong

ILovePuppiesLol (ilovepuppieslol):

would you rather go to the left were nothing is right or go to the right where nothing is left

alones (alones):

@Directrix so it is actually right?? ;3

pooja195 (pooja195):

x^2 + x + 1 = correct

alones (alones):

K \( \color{purple}{Finally~i~got~it~righ} \) thanks pooja<3

Directrix (directrix):

(x^3 + 2x^2 + 2x + 1) ÷ (x + 1) = x^2 + x + 1 = correct My question is this: What is this stuff? z^5+z^4+z^3+21z^2+9z+27z+3 =(z+3)(z^4−2z^3+7z^2+9)/z+3 =z^4−2z^3+7z^2+9 Another problem or what? @AloneS

alones (alones):

@Directrix it's all one problem..that's the long diviosn i wrote out

Directrix (directrix):

Why did you change the x variable to z? I'm not following that. If you understand what's up, then I'll move on. @AloneS

alones (alones):

Wait so I ..no the x and z

alones (alones):

I actually changed z to x but yeah it actually doen't matter ;3

ILovePuppiesLol (ilovepuppieslol):

poopers @pooja195

alones (alones):

@ILovePuppiesLol Be nice to pooja she actually helped not like you

ILovePuppiesLol (ilovepuppieslol):

;~; poopers is my favorite

alones (alones):

Soo. stop spamming Mr.spammer

ILovePuppiesLol (ilovepuppieslol):

ur the spammer </3

ILovePuppiesLol (ilovepuppieslol):

ill race u to 99 ss

alones (alones):

\( \color{White}{ GO~AHEAD} \)